Suppose that Maria spends 84,500 on consumption, her disposable income is $90,000 and her Marginal Propensity to Save is .15. What does Maria's autonomous consumption equal?

Select one:

a. $5,500

b. $6,500

c. $7,000

d. $8,000

e. $9,500.
Answer #1

Ans: $8,000

Explanation:

C = a + bY              [Where, a = Autonomous consumption , b = MPC and Y = Disposable income]

MPC = 1 - MPS = 1 = 0.15 = 0.85

So, 84,500 = a + 0.85(90,000)

84,500 = a + 76,500

a = 84,500 - 76,500 = $8,000